OpenStudy (ksaimouli):

find the equation for a line that is tangent to the graph y=e^x and goes through origin

OpenStudy (anonymous):

argh. I can't remember how to differentiate y = e^x. Do you know?

OpenStudy (anonymous):

\[e^x\]

OpenStudy (anonymous):

Thanks! :)

OpenStudy (ksaimouli):

?

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

\[y=e^xx+0\]

OpenStudy (anonymous):

Is that the equation of the tangent line through the origin? That doesn't seem like it would be a linear equation.

OpenStudy (anonymous):

they meet at \[e^x=e^xx\] \[x=1\] \[y=e\] hence equation is \[y=e\]

OpenStudy (anonymous):

\[y \approx2\]

OpenStudy (anonymous):